OpenStudy (anonymous):

solve 15sec^2theta-20=0

OpenStudy (anonymous):

$$15\sec^2\theta-20=0\\15\sec^2\theta=20\\\sec^2\theta=\frac43\\\sec\theta=\pm\sqrt{\frac43}=\pm\frac2{\sqrt3}\\\cos\theta=\pm\frac{\sqrt3}2\implies\theta=?$$